OpenStudy (anonymous):

Show all work: five more than the square root of a number is 11. what is the number?

jimthompson5910 (jim_thompson5910):

sqrt(x)+5=11 sqrt(x)=11-5 sqrt(x)=6 x=6^2 x=36 So the number is 36